OpenStudy (unklerhaukus):

QUESTION 3 We have seen that the magnetic field at the center of a circular loop of radius \(R\) of current \(I\) is \(B=\frac{\mu_0I}{2R}\). What is a lower bound estimate of the inductance of this single loop of wire?

OpenStudy (vincent-lyon.fr):

Good question. I like it. Do you have any idea?

OpenStudy (unklerhaukus):

i also posted here; but the solution the .sam. got to wasn't even on of the options, http://openstudy.com/users/unklerhaukus#/updates/516680e6e4b0902dbc1d7bb9 i ended up guessing the wrong solution it turns out the answer they wanted is \[\frac{\mu_0\pi R}2\] but i dont really understand can you explain it ?

OpenStudy (vincent-lyon.fr):

B is not uniform on the disc limited by the circuit. It is maximum at the centre. So real flux is less than field times area \(\Phi <B_o \pi R^2=\Large \frac {\mu _o I}{2R}\normalsize \pi R^2=\Large \frac {\mu_o \pi R}{2}\normalsize I\) Since \(\Phi=L I\) , then \(L>\Large \frac {\mu_o \pi R}{2}\)